# Curating art exhibitions using machine learning

### v1

@@ Abstract
Art curatorship (Graham and Cook, 2010) has always been mostly the subjective work of human experts, who, with extensive knowledge of many and diverse artworks, select a few of those to present in communal spaces, spaces that evolved into what we now call art galleries. There are no hard and fast set of rules on how to select these artworks, given a theme which either is presented to the art curator or constructed by her/him. Here we present a series of artificial models – a total of four related models – based on machine learning techniques (a subset of artificial intelligence) that attempt to learn from existing exhibitions which have been curated by human experts, in order to be able to do similar curatorship work. We focus exclusively on the last 25 years of past exhibitions at the Metropolitan Museum of Art in New York, due to the quality of the data available and the physical and time limitations of our research. Our four artificial intelligence models achieve a reasonable ability at imitating these various curators responsible for all those exhibitions, with various degrees of precision and curatorial coherence. In particular, we can conclude two key insights: first, that there is sufficient information in these exhibitions to construct an artificial intelligence model that replicates past exhibitions with an accuracy well above random choices; second, that using feature engineering and carefully designing the architecture of modest size models can make them as good as those using the so-called large language models such as GPT in a brute force approach. We also believe, based on small attempts to use the models in out-of-sample experiments, that given more much more data, it should be possible for these kinds of artificial intelligence agents to be closer and closer to the aesthetic and curatorial judgment of human art curators.
Keywords: art curation, machine learning, embeddings, large language models (LLMs)
@@ 1 Introduction
Artificial intelligence (AI) and its applications are currently a hot topic in academia as well as in the news, particularly with the announcement of the 2024 Nobel Prize in Physics, which was attributed to work by researchers in AI (The Royal Swedish Academy of Sciences, 2024). In the area of art, AI has already been applied in several research fields, including: generating art (images) in any style one wants (Gatys et al., 2016); generating images from pure text (Ramesh et al., 2021; Esser et al., 2021; Radford et al., 2021); translating paintings to possible corresponding real-world images (or vice versa) and other artistic styles (Zhu et al., 2017); discovering similar patterns in disparate paintings (Saleh et al., 2016; Shen et al., 2019); clustering similar artworks (Hamilton et al., 2021; Castellano and Vessio, 2022); style/artistic school and artist identification/classification (Keren, 2002; Shamir et al., 2010; Shamir and Tarakhovsky, 2012; Cetinic and Grgic, 2013; Saleh and Elgammal, 2015; The Museum of Modern Art, 2025b); painting to text/semantic knowledge extraction (Vinyals et al., 2015); aesthetic evaluation and artistic opinion extraction (Cetinic et al., 2019); and museum/exhibition curation and artwork selection/recommendation (He et al., 2016; Krysa and Moscoso, 2019; Messina et al., 2019, 2020; Bowen et al., 2020; Fosset et al., 2022; Yilma and Leiva, 2023; Nasher Museum of Art at Duke University, 2023; Schaerf et al., 2023; Herman and Moruzzi, 2024; von Davier et al., 2024; Srinivasan, 2024; Gordon, 2025).
@@
Regarding the last field of research, museum and exhibition curation using AI, most of the work seems to focus on either finding similar artworks to a given (seed) artwork, or using a prompt (e.g., for the OpenAI GPT-4 (OpenAI, 2023)), to select appropriate artworks. Here we propose a different and modest approach that does not seem to have been explored yet. We propose to use existing collections of actual past art exhibitions (see e.g. ArtLAS Project, 2025; University of Vienna, 2025; Museum of Modern Art (2025), MoMA; National Gallery of Art, 2025; The Metropolitan Museum of Art, 2025a, and others), to create a recommender system (for an example of recommender systems see Hu et al., 2008) using AI that imitates what human curators do. That is, we propose to teach AI to learn from past curatorship. The closest of the above-mentioned works to our proposal is the one by the Nasher Museum of Art at Duke University (Nasher Museum of Art at Duke University, 2023), which used OpenAI’s ChatGPT platform to curate exhibitions. The main difference with respect to our work here is that our modelling is absolutely restricted to a given set of artworks (therefore no hallucinations giving artworks outside that set), and more importantly and crucially, we have made the models learn from past exhibitions using those artworks and only learn from that set of information.
@@
This is how we propose to train our models. The training set would consist of the input pairs $(x,y)$, where $x$ is the embedding vector (see e.g. Kenton and Toutanova, 2019, on text-to-vector embeddings) of the exhibition title and main text, while $y$ is made of the features extracted from the images of the artworks exhibited. Optionally, one could add to $y$ the embedding vector of the caption of artworks that were selected by the human curators, if available. [^1] We could consider also using the information on who was the human curator, for a more nuanced system, that could recognise and create an exhibition on the style of a certain human curator, but given the scarcity of information available, that is wildly optimistic. After the training phase, users could input on the test phase, a title and a description of a hypothetical exhibition, and the system would choose the appropriate artworks to include, and would choose only from the given set of artworks.
@@ 2 Dataset
Although there are several datasets available (see e.g. ArtLAS Project, 2025; University of Vienna, 2025; Museum of Modern Art (2025), MoMA; National Gallery of Art, 2025; The Metropolitan Museum of Art, 2025a, and others), we have decided to restrict ourselves to the Metropolitan Museum of Art in New York (hereafter Met Museum), as it had two essential ingredients for our intended research project, namely: an extensive and well formatted dataset on past exhibitions and their works of art, and the corresponding freely available dataset on details and tags associated with those works of art, i.e., their metadata. Furthermore, the merging of databases could have proven a very complex task, as these sets all have different metadata associated to the artworks, and the data format come in a large variety.
On the first part of this dataset, we have focused on the data from the Met Museum’s website (The Metropolitan Museum of Art, 2025b), and extracted directly the full set of archived exhibitions and where available, its corresponding works of art (or rather the object id, an integer that clearly identifies the artwork within the Met Museum’s associated artwork database). This dataset included a total of $1636$ exhibitions from 2000 to 2025 inclusive, although these included duplicated exhibitions, as some of those encompass several years. Also, some exhibitions did not contain the list of object ids to map to works of art and/or only contained objects temporarily loaned to the Met Museum and therefore not in the database of objects. After removing those and other exhibitions we could not use, we end up with a much lower value of $236$ exhibitions. We depict in Table 1 a summary of that data, which corresponds to a download at the date of 24 February 2025 (naturally, the Met Museum may have released more webpages with newer exhibitions, but we have frozen our data at this date).
@@
Table 1: Exhibition statistics by year for the downloaded exhibitions from the Metropolitan Museum of Art of New York, which corresponds to a download at the date of 24 February 2025. It shows the total number of exhibitions per year, the number of those that have the artworks explicitly stated (using Object IDs that match the artwork database from the Met Museum) and the number of artworks found. As some exhibitions run across year ends, we also show the corresponding total all years’ value of the number of exhibitions that have known Object IDs and the corresponding lower number of actual artworks in that unique set of exhibitions. Furthermore, we also show those numbers when one removes exhibitions that encompass only artworks not on the Met Museum’s database, i.e., exhibitions that are made of loaned artworks. The actual usable number of exhibitions for our research project is therefore only 236, and the number of artworks a total of $10,470$ (with possible repeats across exhibitions). The number of actual unique artworks is $9289$. Finally, we show the number of words in each year for all the titles and descriptions for all exhibitions within that year, and the total number of words for all exhibitions with and without duplication, and for the actual exhibitions we shall use for our modelling. For comparison, it is speculated that the OpenAI’s gpt-4o-mini model uses several billion words for its training.
The exhibition dataset, consisting of a grand total of 236 exhibitions after removal of multi-year duplicates and made of exhibitions for which we have object ids in the corresponding Met Museum’s artworks object id to metadata file, has the following format, as depicted in Listing 1. For each exhibition, we have the title of the exhibition and the description of the exhibition (called overview_text), all in English and with the replacement of any diacritics (accents, umlauts, tildes, etc.) and non-ASCII characters, and then we have a variable number of object_ids, which identify the artworks within the Met Museum.222 Note that this id is not a worldwide id, but an internal Met Museum. There are some attempts to catalogue and identify all artworks with a unique worldwide id, e.g.,: Wikidata Qxxxxx ids — “Les Demoiselles d’Avignon” by Pablo Picasso has a unique id Q910199; the International Image Interoperability Framework (IIIF Consortium, 2024); the Europeana project (Europeana Foundation, 2024); among many others. For each object_id, we have a series of metadata elements, which can be a single string, a list of strings or an empty list with no element in it.
@@
Listing 1: The excerpt of the raw format of our data, showing an exhibition by the Met Museum, from the year 2000 (https://www.metmuseum.org/exhibitions/listings/2000/spanish-renaissance), displaying the title and overview text describing the purpose and essence of the exhibition, and some of the mappings to the actual object_ids and their respective metadata from The Metropolitan Museum of Art (2025a).
@@
On the second part of this dataset, we used the Met Museum’s Open Access CSV in The Metropolitan Museum of Art (2025a) data, which contains the museum’s entire collection index of works of art (a total of $484,956$ works of art) and its associated metadata. The majority of these objects did not show up on any exhibition on the list of exhibitions that we describe above. The metadata consists of many fields (54 in total), but we focus only on “Department”, “Object Name”, “Title”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”, together with “Object ID” to link to the exhibitions’ extract. Note that later on, for the machine learning approaches we ignore “Object Name”, “Title”, as these are somehow too arbitrary to be used as meaningful metadata for our modelling. The count of the top strings in each of the fields we consider (“Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) is depicted in Table 2. We shall call these set of all these fields generalised tags (as opposed to only tags, which is already used by the Met Museum as a column name itself in its dataset of all artworks.) There are a total of $69,566$ generalised tag strings in our full dataset of 236 exhibitions, of which $8591$ are unique generalised tags strings. Below we display a very small subset of this dataset in Table LABEL:tab:art_objects as an example for the reader.
Table 3: Excerpt of the Met Museum database of all its $484,956$ artworks. We select only a few metadata fields, which we call generalised tags. The total number of fields present on the database is 54 in total, but we use only “Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”, together with “Object ID” to link to the exhibitions’ extract. Above we display “Object Name”, and “Title” for information purposes only, these are not used in the modelling. We truncate some of the values of the metadata fields with …for presentation only in this table, we use the full strings in our modelling code. For some of the values we have more than one string, e.g., “Female Nudes|Bathing”, which means we have a list of strings, two strings in this particular case. The “|” or pipe symbol is used in comma separated files (csv files) as the one given by the Met Museum in its Application Programming Interface (API) (The Metropolitan Museum of Art, 2025a) to expose its database to external users.
@@
Field Non-empty Entries
Department $484,956$
Object Name $484,956$
Title $484,956$
Artist Display Name $282,513$
Object Begin Date $484,956$
Medium $484,950$
Classification $406,239$
Tags $192,455$
All Fields Non-empty $120,713$
Table 4: Non-empty Fields in Metropolitan Museum of Art Dataset. Notice that the fields “Object Name”, “Title” are used for information purposes only, and do not enter the machine learning approaches.
@@ 3 Methodologies and results
We attempt this mapping between inputs and outputs via four distinct approaches. All our approaches draw on methods developed in the field of Natural Language Processing (NLP) – see e.g., Young et al. (2018) for a review and references therein. Schematically, all our approaches can be summarised by Figure 1, which displays a toy model where we input an exhibition title and description, and output either a list of generalised tags (which is then mapped to the nearest objects that match those tags), or an actual direct list of artworks. Naturally, the neural networks we use are much larger, both in depth and width. Furthermore, some of those neural networks can be much more complex, incorporating other types of connections including loops.
@@
First, we use just the data as it is, without external help, by using its internal statistics. That is to say, we use the probability of finding a word in English within the corpus [^3] A corpus is a collection of texts that are used for linguistic/statistic analysis. of exhibition titles and descriptions. In other words, we create a model to represent those words within the language patterns of that corpus. In NLP, this is usually done by using the so-called text vectorisation, which computes word (more accurately, it uses tokens, which can be words, subsets of words or letters) frequency and relative distribution within the corpus (Salton et al., 1975). Therefore, we create a text vectorisation based on the contents of all the titles concatenated with the overview (description) texts for all exhibitions taken together. That is to say, we transform the input texts into numeric vectors using the self-contained statistics of the exhibition titles and texts (see e.g., Mikolov et al., 2013, on embeddings). Second, we now use external help, in the form of the state-of-the-art OpenAI embedding model text-embedding-3-large that can perform very high-quality text vectorisation. In this sense, we are using the statistical analysis performed by OpenAI on their (much) larger corpus of texts to determine the best (numeric) vectors that represent strings of text. These models also have the advantage of capturing semantic similarity in a metric way (see e.g., Corley and Mihalcea, 2005, for details of semantic similarity in NLP), i.e., they convert textual data into numeric vectors where similar words/sentences/paragraphs have similar vector representation, being near, in the high-dimensional space [^4] We note here that the dimension of these spaces can be very large, typically of the order of thousands or even tens of thousands of dimensions, making it impossible to clearly visualise these vector representations. characterising the texts. In our first model, we do not capture semantic similarity, only word frequency, so this second approach, has, in principle, some advantages. Third, we use OpenAI embedding models to transform both the input, the title and description of the exhibitions, and the output, for which we use the concatenated strings made of all the metadata associated with all the artworks for each of one the exhibitions. We then can find the nearest embedding vector for an input text and from that find the nearest artworks’ metadata embedding vectors for that text, giving us the nearest artworks. Fourth, and finally, we use the full power and knowledge encapsulated in the OpenAI models, by using the OpenAI Application Programming Interface (API [^5] An API is a collection or set of rules and specifications that allows to call code or databases remotely without having to store or host the entire infrastructure locally.) to call OpenAI models to ask for a direct and full list of generalised tags suggested for a given exhibition description (title concatenated with the overview text). We therefore train one of the latest OpenAI models, “gpt-4o-mini”, using a technique called fine-tuning (Devlin et al., 2019; Yosinski et al., 2014). Fine-tuning models is a common approach in the field of neural networks, taking an already pre-trained network and improving it (or fine-tuning it) to a more specific domain (see e.g., Covas, 2020, 2023, for two examples in astrophysics and in finance). After fine-tuning, we ask the resulting model via the API for lists of generalised tags (without forcing the number of predicted artworks) given an input text (which the fine-tuned model will decided how to represent and embed numerically).
@@
For all approaches, we divide our dataset in Table 1 into a training and validation sets, as it is standard in machine learning approaches. We do this in the proportion $80\%$ examples for training (a total of 188 examples) and $20\%$ samples for validation or in-sample testing (a total of 48 examples), taking the total to the above-mentioned global set of 236 exhibitions. [^8] The validation set is a subset of examples drawn from the same sample – this means it presumably follows the same probability distribution function – but is not used for training, just for checking accuracy. On the other hand, test sets can be in-sample (from the same probability distribution) or out-of-sample (from possibly a different distribution), and the latter are used to test the ability of the model to actually generalise. We train the models on the training set, [^9] We train and optimise the neural network using the so-called Adam optimiser (Kingma, 2014), which is implemented within Google’s TensorFlow deep learning framework (Abadi et al., 2016) in our Python code. and then test our ability to predict the appropriate metadata (“Department”,“Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) and ultimately the artworks for the input text in the validation set, comparing the predicted versus actual metadata or generalised tags and the list of predicted versus actual artworks. All our training optimises to minimise the error difference between the actual output (on the training set) and the predicted output. We also calculate the error difference within the validation set to monitor over-fitting, a common problem with machine learning models (Geman et al., 1992). Only after training the model, and checking the results for the training and validation sets, we then test out-of-sample on a made-up test set, where we create an input hypothetical exhibition title and description, in the hope that our model can select appropriate artworks for it.

### v2

@@ Abstract
Here we present a series of artificial models – a total of four related models – based on machine learning techniques that attempt to learn from existing exhibitions which have been curated by human experts, in order to be able to do similar curatorship work. Out of our four artificial intelligence models, three achieve a reasonable ability at imitating these various curators responsible for all those exhibitions, with various degrees of precision and curatorial coherence. In particular, we can conclude two key insights: first, that there is sufficient information in these exhibitions to construct an artificial intelligence model that replicates past exhibitions with an accuracy well above random choices; and second, that using feature engineering and carefully designing the architecture of modest size models can make them almost as good as those using the so-called large language models such as GPT in a brute force approach.
Keywords: art curation; machine learning; embeddings; large language models (LLMs)
@@ 1 Introduction
Artificial intelligence (AI) and its applications are currently a hot topic in academia as well as in the news, particularly with the announcement of the 2024 Nobel Prize in Physics, which was attributed to work by researchers in AI [The Royal Swedish Academy of Sciences, 2024]. In the area of art, AI has already been applied in several research fields, including: generating art (images) in any style one wants [Gatys et al., 2016]; generating images from pure text [Ramesh et al., 2021, Esser et al., 2021, Radford et al., 2021]; translating paintings to possible corresponding real-world images (or vice versa) and other artistic styles [Zhu et al., 2017]; discovering similar patterns in disparate paintings [Saleh et al., 2016, Shen et al., 2019]; clustering similar artworks [Hamilton et al., 2021, Castellano and Vessio, 2022]; style/artistic school and artist identification/classification [Keren, 2002, Shamir et al., 2010, Shamir and Tarakhovsky, 2012, Cetinic and Grgic, 2013, Saleh and Elgammal, 2015, The Museum of Modern Art, 2025b]; painting to text/semantic knowledge extraction [Vinyals et al., 2015]; aesthetic evaluation and artistic opinion extraction [Cetinic et al., 2019]; and museum/exhibition curation and artwork selection/recommendation [He et al., 2016, Krysa and Moscoso, 2019, Messina et al., 2019, 2020, Bowen et al., 2020, Fosset et al., 2022, Yilma and Leiva, 2023, Nasher Museum of Art at Duke University, 2023, Schaerf et al., 2023, Herman and Moruzzi, 2024, von Davier et al., 2024, Srinivasan, 2024, Gordon, 2025, Li, 2025].
@@
Regarding the last field of research, museum and exhibition curation using AI, most of the work seems to focus on either finding similar artworks to a given (seed) artwork, or using a prompt (e.g., for the OpenAI GPT-4 [OpenAI, 2023]), to select appropriate artworks. Here we propose a different and modest approach that does not seem to have been explored yet. We propose to use existing collections of actual past art exhibitions [see e.g., ArtLAS Project, 2025, University of Vienna, 2025, Museum of Modern Art (2025), MoMA, National Gallery of Art, 2025, The Metropolitan Museum of Art, 2025a, and others], to create a recommender system [for an example of recommender systems see Hu et al., 2008] using AI that imitates what human curators do. That is, we propose to teach AI to learn from past curatorship. The closest of the above-mentioned works to our proposal is the one by the Nasher Museum of Art at Duke University [Nasher Museum of Art at Duke University, 2023], which used OpenAI’s ChatGPT platform to curate exhibitions. The main difference with respect to our work here is that our modelling is absolutely restricted to a given set of artworks (therefore no hallucinations giving artworks outside that set), and more importantly and crucially, we have made the models learn from past exhibitions using those artworks and only learn from that set of information.
@@
This is how we propose to train our models. The training set would consist of the input pairs $(x,y)$, where $x$ is the embedding vector [see e.g., Kenton and Toutanova, 2019, on text-to-vector embeddings] of the exhibition title and main text, while $y$ is made of the features extracted from the images of the artworks exhibited. Optionally, one could add to $y$ the embedding vector of the caption of artworks that were selected by the human curators, if available. [^1] We could consider also using the information on who was the human curator, for a more nuanced system, that could recognise and create an exhibition on the style of a certain human curator, but given the scarcity of information available, that is wildly optimistic. After the training phase, users could input on the test phase, a title and a description of a hypothetical exhibition, and the system would choose the appropriate artworks to include, and would choose only from the given set of artworks.
@@ 2 Dataset
Although there are several datasets available [see e.g., ArtLAS Project, 2025, University of Vienna, 2025, Museum of Modern Art (2025), MoMA, National Gallery of Art, 2025, The Metropolitan Museum of Art, 2025a, and others], we have decided to restrict ourselves to the Metropolitan Museum of Art in New York (hereafter Met Museum), as it had two essential ingredients for our intended research project, namely: an extensive and well formatted dataset on past exhibitions and their works of art, and the corresponding freely available dataset on details and tags associated with those works of art, i.e., their metadata. Furthermore, the merging of databases could have proven a very complex task, as these sets all have different metadata associated to the artworks, and the data formats come in a large variety.
On the first part of this dataset, we have focused on the data from the Met Museum’s website [The Metropolitan Museum of Art, 2025b], and extracted directly the full set of archived exhibitions and where available, its corresponding works of art (or rather the object id, an integer that clearly identifies the artwork within the Met Museum’s associated artwork database). This dataset included a total of $1636$ exhibitions from 2000 to 2025 inclusive, although these included duplicated exhibitions, as some of those encompass several years. Also, some exhibitions did not contain the list of object ids to map to works of art and/or only contained objects temporarily loaned to the Met Museum and therefore not in the database of objects. After removing those and other exhibitions we could not use, we end up with a much lower value of $236$ exhibitions. We depict in Table 1 a summary of that data, which corresponds to a download on the date of 24 February 2025 (naturally, the Met Museum may have released more webpages with newer exhibitions, but we have frozen our data on this date).
@@
Table 1: Exhibition statistics per year for the downloaded exhibitions from the Metropolitan Museum of Art of New York, which corresponds to a download on the date of 24 February 2025. It shows the total number of exhibitions per year, the number of those that have the artworks explicitly stated (using Object IDs that match the artwork database from the Met Museum) and the number of artworks found. As some exhibitions run across year ends, we also show the corresponding total all years’ value of the number of exhibitions that have known Object IDs and the corresponding lower number of actual artworks in that unique set of exhibitions. Furthermore, we also show those numbers when one removes exhibitions that encompass only artworks not on the Met Museum’s database, i.e., exhibitions that are made of loaned artworks. The actual usable number of exhibitions for our research project is therefore only 236, and the number of artworks a total of $10,470$ (with possible repeats across exhibitions). The number of actual unique artworks is $9289$. Finally, we show the number of words in each year for all the titles and descriptions for all exhibitions within that year, and the total number of words for all exhibitions with and without duplication, and for the actual exhibitions we shall use for our modelling. For comparison, it is speculated that the OpenAI’s gpt-4o-mini model uses several billion words for its training.
The exhibition dataset, consisting of a grand total of 236 exhibitions after removal of multi-year duplicates and made of exhibitions for which we have object ids in the corresponding Met Museum’s artworks object id to metadata file, has the following format, as depicted in Listing 1. For each exhibition, we have the title of the exhibition and the description of the exhibition (called overview_text), all in English and with the replacement of any diacritics (accents, umlauts, tildes, etc.) and non-ASCII characters, and then we have a variable number of object_ids, which identify the artworks within the Met Museum. [^2] Note that this id is not a worldwide id, but an internal Met Museum id. There are some attempts to catalogue and identify all artworks with a unique worldwide id, e.g.,: Wikidata Qxxxxx ids — “Les Demoiselles d’Avignon” by Pablo Picasso has a unique id Q910199; the International Image Interoperability Framework [IIIF Consortium, 2024]; the Europeana project [Europeana Foundation, 2024]; among many others. For each object_id, we have a series of metadata elements, which can be a single string, a list of strings or an empty list with no elements in it.
@@
Listing 1: The excerpt of the raw format of our data, showing an exhibition by the Met Museum, from the year 2000 (https://www.metmuseum.org/exhibitions/listings/2000/spanish-renaissance), displaying the title and overview text describing the purpose and essence of the exhibition, and some of the mappings to the actual object_ids and their respective metadata from The Metropolitan Museum of Art [2025a].
On the second part of this dataset, we used the Met Museum’s Open Access CSV in The Metropolitan Museum of Art [2025a] data, which contains the museum’s entire collection index of works of art (a total of $484,956$ works of art) and its associated metadata. The majority of these objects did not show up on any exhibition on the list of exhibitions that we describe above. The metadata consists of many fields (54 in total), but we focus only on “Department”, “Object Name”, “Title”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”, together with “Object ID” to link to the exhibitions’ extract. Note that later on, for the machine learning approaches we ignore “Object Name”, “Title”, as these are somehow too arbitrary to be used as meaningful metadata for our modelling. The count of the top strings in each of the fields we consider (“Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) is depicted in Table 2. We shall call these set of all these fields generalised tags (as opposed to only tags, a name which is already used by the Met Museum as a column name in its dataset of all artworks.) There are a total of $69,566$ generalised tag strings in our full dataset of 236 exhibitions, of which $8591$ are unique generalised tags strings. Below we display a very small subset of this dataset in Table 4 as an example for the reader.
@@
Field Non-empty Entries
Department $484,956$
Object Name $484,956$
Title $484,956$
Artist Display Name $282,513$
Object Begin Date $484,956$
Medium $484,950$
Classification $406,239$
Tags $192,455$
All Fields Non-empty $120,713$
Table 3: Non-empty Fields in Metropolitan Museum of Art Dataset. Notice that the fields “Object Name” and “Title” are used for information purposes only, and do not enter the machine learning approaches.
@@
Table 4: Excerpt of the Met Museum database of all its $484,956$ artworks. We select only a few metadata fields, which we call generalised tags. The total number of fields present on the database is 54 in total, but we use only “Department”, “Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”, together with “Object ID” to link to the exhibitions’ extract. Above we display “Object Name”, and “Title” for information purposes only, these are not used in the modelling. We truncate some of the values of the metadata fields with “…” for presentation only in this table, we use the full strings in our modelling code. For some of the values we have more than one string, e.g., “Female Nudes|Bathing”, which means we have a list of strings, two strings in this particular case. The “|” or pipe symbol is used in comma separated files (csv files) as the one given by the Met Museum in its Application Programming Interface (API) [The Metropolitan Museum of Art, 2025a] to expose its database to external users.
@@ 3 Methodologies and results
We attempt this mapping between inputs and outputs via four distinct approaches. All our approaches draw on methods developed in the field of Natural Language Processing (NLP) – see e.g., Young et al. [2018] for a review and references therein. Schematically, all our approaches can be summarised by Figure 1, which displays a toy model where we input an exhibition title and description, and output either a list of generalised tags (which is then mapped to the nearest objects that match those tags), or an actual direct list of artworks. Naturally, the neural networks we use are much larger, both in depth and width. Furthermore, some of those neural networks can be much more complex, incorporating other types of connections including loops.
@@
First, we use just the data as it is, without external help, by using its internal statistics. That is to say, we use the probability of finding a word in English within the corpus [^3] A corpus is a collection of texts that are used for linguistic/statistical analysis. of exhibition titles and descriptions. In other words, we create a model to represent those words within the language patterns of that corpus. In NLP, this is usually done by using the so-called text vectorisation, which computes word (more accurately, it uses tokens, which can be words, subsets of words or letters) frequency and relative distribution within the corpus [Salton et al., 1975]. Therefore, we create a text vectorisation based on the contents of all the titles concatenated with the overview (description) texts for all exhibitions taken together. That is to say, we transform the input texts into numeric vectors using the self-contained statistics of the exhibition titles and texts [see e.g., Mikolov et al., 2013, on embeddings]. Second, we now use external help, in the form of the state-of-the-art OpenAI embedding model text-embedding-3-large that can perform very high-quality text vectorisation. In this sense, we are using the statistical analysis performed by OpenAI on their (much) larger corpus of texts to determine the best (numeric) vectors that represent strings of text. These models also have the advantage of capturing semantic similarity in a metric way [see e.g., Corley and Mihalcea, 2005, for details of semantic similarity in NLP], i.e., they convert textual data into numeric vectors where similar words/sentences/paragraphs have similar vector representation, being near, in the high-dimensional space [^4] We note here that the dimension of these spaces can be very large, typically of the order of thousands or even tens of thousands of dimensions, making it impossible to clearly visualise these vector representations. characterising the texts. In our first model, we do not capture semantic similarity, only word frequency, so this second approach, has, in principle, some advantages. Third, we use OpenAI embedding models to transform both the input, the title and description of the exhibitions, and the output, for which we use the concatenated strings made of all the metadata associated with all the artworks for each of one the exhibitions. We then can find the nearest embedding vector for an input text and from that find the nearest artworks’ metadata embedding vectors for that text, giving us the nearest artworks. Fourth, and finally, we use the full power and knowledge encapsulated in the OpenAI models, by using the OpenAI Application Programming Interface (API [^5] An API is a collection or set of rules and specifications that allows to call code or databases remotely without having to store or host the entire infrastructure locally.) to call OpenAI models to ask for a direct and full list of generalised tags suggested for a given exhibition description (title concatenated with the overview text). We therefore train one of the latest OpenAI models, “gpt-4o-mini”, using a technique called fine-tuning [Devlin et al., 2019, Yosinski et al., 2014]. Fine-tuning models is a common approach in the field of neural networks, taking an already pre-trained network and improving it (or fine-tuning it) to a more specific domain [see e.g., Covas, 2020, 2023, for two examples in astrophysics and in finance]. After fine-tuning, we ask the resulting model via the API for lists of generalised tags (without forcing the number of predicted artworks) given an input text (which the fine-tuned model will decided how to represent and embed numerically).
@@
For all approaches, we divide our dataset in Table 1 into a training and validation sets, as it is standard in machine learning approaches. We do this in the proportion $80\%$ examples for training (a total of 188 examples) and $20\%$ samples for validation or in-sample testing (a total of 48 examples), taking the total to the above-mentioned global set of 236 exhibitions. [^8] The validation set is a subset of examples drawn from the same sample – this means it presumably follows the same probability distribution function – but is not used for training, just for checking accuracy. On the other hand, test sets can be in-sample (from the same probability distribution) or out-of-sample (from possibly a different distribution), and the latter are used to test the ability of the model to actually generalise. We train the models on the training set, [^9] We train and optimise the neural network using the so-called Adam optimiser [Kingma, 2014], which is implemented within Google’s TensorFlow deep learning framework [Abadi et al., 2016] in our Python code. and then test our ability to predict the appropriate metadata (“Department”,“Artist Display Name”, “Object Begin Date”, “Medium”, “Classification”, “Tags”) and ultimately the artworks for the input text in the validation set, comparing the predicted versus actual metadata or generalised tags and the list of predicted versus actual artworks. All our training optimises to minimise the error difference between the actual output (on the training set) and the predicted output. We also calculate the error difference within the validation set to monitor over-fitting, a common problem with machine learning models [Geman et al., 1992]. Only after training the model, and checking the results for the training and validation sets, we then test out-of-sample on a made-up test set, where we create an input hypothetical exhibition title and description, in the hope that our model can select appropriate artworks for it.